OpenStudy (anonymous):

49 divided by 7over13

OpenStudy (espex):

Is it \[\frac{(\frac{49}{7})}{13}\]or \[\frac{49}{(\frac{7}{13})}\]

OpenStudy (anonymous):

THE SECOND ONE 49 DIVIDE 7OVER13

OpenStudy (anonymous):

It equals 49 times the reciprocal of 7 over 13. so 49 times 13 over 7.

OpenStudy (espex):

As @c0rtez pointed out, when dividing fractions you are able to simply multiply the reciprocal. So: \[\frac{49}{(\frac{7}{13})}\] becomes \[49 * \frac{13}{7}\]

OpenStudy (anonymous):

remember this when it comes to dividing fractions, in your case its \[49 \div \frac{ 7 }{ 13 }\] leave it switch it flip it Ok, so what this means is that you leave 49 the way it is (LEAVE IT). you switch the divided by sign to a multiplications sign (SWITCH IT). and you flip the 7 and 13 to 13 over 7 (FLIP IT). after this you will end up with: \[49 \div \frac{ 13 }{7 }\]